OpenStudy (anonymous):

Using the definition of derivative, and find f'(x). Then find f'(-2), f'(0), and f'(3) when the derivative exists.

OpenStudy (anonymous):

What is \(f(x)\)?

OpenStudy (anonymous):

its number 11 from the attachment

OpenStudy (anonymous):

f(x)=3x-7

OpenStudy (anonymous):

\[f'(x)=\lim_{h\to0}\frac{\bigg(3(x+h)-7\bigg)-\bigg(3x-7\bigg)}{h}\] \[f'(x)=\lim_{h\to0}\frac{3x+3h-7-3x+7}{h}\] \[f'(x)=\lim_{h\to0}\frac{3h}{h}\] \(\vdots\)

OpenStudy (anonymous):

\[f \prime \left( x \right)=\lim_{h \rightarrow 0}\frac{ f \left( x+h \right)-f \left( x \right) }{ h }\]

OpenStudy (anonymous):

and what do i do with the f'(-2), f'(0),f'(3)??

OpenStudy (anonymous):

Do i plug it in for x

OpenStudy (anonymous):

as calculated above f'(x)=3 now you can find ?

OpenStudy (anonymous):

so i do plug it in?

OpenStudy (anonymous):

whatever you plug in the value will come out to be 3.

OpenStudy (anonymous):

can you do me the first problem so i can see visually and understand it because im lost

OpenStudy (anonymous):

f'(x)=3 f'(-2)=3 f(0)=3 ..........

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

there is no x in f'(x) hence it will remain 3